OpenStudy (anonymous):

i have 3 qestions, i will post in the comments

OpenStudy (anonymous):

first: \[\sqrt{5}(8+3\sqrt{6}\]

OpenStudy (anonymous):

\[8\sqrt{5}+3\sqrt{30}\]

OpenStudy (anonymous):

distributive property

OpenStudy (anonymous):

second: \[(7+\sqrt{2})(8-3\sqrt{2})\]

OpenStudy (anonymous):

third: \[(8+\sqrt{5})(8-\sqrt{5})\]

OpenStudy (anonymous):

and identify the conjugate of \[4+\sqrt{7}\]

OpenStudy (anonymous):

conjugate of \[a+\sqrt{b}\] is \[a-\sqrt{b}\]

OpenStudy (anonymous):

\[(8+\sqrt{5})(8-\sqrt{5})\] \[(a+b)(a-b)=a^2-b^2\] so you get \[(8+\sqrt{5})(8-\sqrt{5})=8^2-\sqrt{5}^2=64-5\]

OpenStudy (anonymous):

\[(7+\sqrt{2})(8-3\sqrt{2})\] like \[(7+x)(8-3x)\] multiply out and see what you get